OpenStudy (anonymous):

I will give you a medal if can help me Solve the system using the elimination method. 2x+4y=-10 -3x-2y=-4 Please show how you got it :)

OpenStudy (anonymous):

2x+4y=-10 -3x-2y=-4 -2y = 3x - 4 y = (-3/2)x + 2 2x + 4((-3/2)x + 2) = -10 x = 9/2 Substitute back in to solve for y. (9/2, -19/4)

OpenStudy (anonymous):

Thanks